OpenStudy (anonymous):

Solve the equation. Check the solution. -2/(x+4)=4/(x+3)

OpenStudy (anonymous):

A. -13/6 B. -11 C. -8/3 D. -11/3

OpenStudy (superdavesuper):

start by multiplying both left n right hand side by (x+4)*(x+3) first: [-2/(x+4)]*(x+4)*(x+3)=[4/(x+3)]*(x+4)*(x+3) what will u get?

OpenStudy (anonymous):

they cancel each other out, right?

OpenStudy (anonymous):

@agent0smith

OpenStudy (superdavesuper):

Yes, so how will the eqn look like after the canceling each other out?

OpenStudy (anonymous):

0=0?

OpenStudy (superdavesuper):

Nope...I will do the left hand side for u: [-2/(x+4)]*(x+4)*(x+3) =-2*(x+3)=-2x-6 That will be =[4/(x+3)]*(x+4)*(x+3) So would you cancel n simplify that plz?

OpenStudy (agent0smith):

Unfortunately the equation editor still doesn't work. Makes this so much harder than it should be.

OpenStudy (anonymous):

4/x+4?

OpenStudy (anonymous):

because the x+3's cancel??

OpenStudy (agent0smith):

^^^ 4*(x+4)

OpenStudy (anonymous):

ahh okay, so then what?

OpenStudy (superdavesuper):

@agent0smith i know >=< @HelennnnnnC the left and right hand side are equal so: -2x-6=4x+4 u should be able to solve x from here :)

OpenStudy (agent0smith):

@superdavesuper i was just correcting helen's mistake, not yours :)

OpenStudy (anonymous):

would it be 6x=-10

OpenStudy (anonymous):

so -10/6 = -5/3?

OpenStudy (anonymous):

but thats not one of the answers

OpenStudy (agent0smith):

4*(x+4) = 4x+16, i think you missed that helen -2x-6=4x+16

OpenStudy (anonymous):

oh so it's 6x=-22

OpenStudy (superdavesuper):

Ooops...me bad in my last post - plz fix MY mistakes as well @agent0smith :) keep going @HelennnnnnC u are almost there!

OpenStudy (anonymous):

so then it's -11/3?

OpenStudy (superdavesuper):

that is correct! :)